Find the characteristic values and characteristic functions of the storm lionville problem

d2y/dx2+λy=0, y(0)=0,y(π)=0,

where λ is negative

1
Expert's answer
2021-10-25T14:50:05-0400

if "\lambda<0" then "\lambda=-\mu^2", where µ is real and non-zero. The general solution:

"y(x)=Ae^{\mu x}+Be^{-\mu x}"


for boundary conditions:

"y(0)=A+B=0\implies A=-B"

"y(\pi)=Ae^{\mu \pi}+Be^{-\mu \pi}=0\implies \mu=0"


So, there is only trivial solution with A=0,B=0:

"y(x)=0"